Elmer had a job pulling weeds that paid $7 an hour and he worked at most 42 hours a week. If x is the hours he worked in a week and y is money he made in a week, write the domain and range of the function as compound inequalities.
1
Expert's answer
2020-12-15T01:18:09-0500

"0\le X\le 42"

"0\le Y \le 294"